OpenStudy (anonymous):

When asked to factor the trinomial 9x^2+12x+4 , a student gives the answer (3x - 2)(3x - 2). What is one thing wrong with this answer? A. 3 is also a factor of this trinomial B. There is nothing wrong with the answer C. The factors are not simplified D. The minus signs should be plus signs

OpenStudy (aum):

9x^2+12x+4 = 9x^2 + 6x + 6x + 4 = 3x(3x + 2) + 2(3x + 2) = (3x + 2)(3x + 2)

OpenStudy (anonymous):

so D